What concentration of sulfuric acid will make a solution of pH=3.4
1
Expert's answer
2020-09-29T06:51:20-0400

We know

pH=-log(H+)

or (H+)=10^-pH=10^-3.4=0.00039

Formula of sulfuric acid is H2SO4

Each molecule contain 2 H

Concentration is 0.00039/2=0.00019(M)